OpenStudy (anonymous):

Simplify the expression. 216^5/6 I keep getting 2^5/2 * 3^5/2 but these are the only answer choices. 6 36 sqrt 6 6^5sqrt6 180

OpenStudy (triciaal):

I got C what are the factors of 216= 2*2*2*3*3*3 can be expressed as 6*6*6 = 6^3 when you raise a power to a power multiply the powers and sqrt = power 1/2 apply to this problem 216^5/6 = 6^3*(5/6) = 6^5/2 = 6^5sqrt6

OpenStudy (triciaal):

by the way your answer is correct a^n *b^n = (ab)^n you have 2^5/2 * 3^5/2 = (2*3)^5/2 = 6^5/2

OpenStudy (anonymous):

Thank you so much!